# Training the Untrainable: Introducing Inductive Bias via Representational Alignment

### v1

@@ Abstract
We demonstrate that architectures which traditionally are considered to be ill-suited for a task can be trained using inductive biases from another architecture. Networks are considered untrainable when they overfit, underfit, or converge to poor results even when tuning their hyperparameters. For example, plain fully connected networks overfit on object recognition while deep convolutional networks without residual connections underfit. The traditional answer is to change the architecture to impose some inductive bias, although what that bias is remains unknown. We introduce guidance, where a guide network guides a target network using a neural distance function. The target is optimized to perform well and to match its internal representations, layer-by-layer, to those of the guide; the guide is unchanged. If the guide is trained, this transfers over part of the architectural prior and knowledge of the guide to the target. If the guide is untrained, this transfers over only part of the architectural prior of the guide. In this manner, we can investigate what kinds of priors different architectures place on untrainable networks such as fully connected networks. We demonstrate that this method overcomes the immediate overfitting of fully connected networks on vision tasks, makes plain CNNs competitive to ResNets, closes much of the gap between plain vanilla RNNs and Transformers, and can even help Transformers learn tasks which RNNs can perform more easily. We also discover evidence that better initializations of fully connected networks likely exist to avoid overfitting. Our method provides a mathematical tool to investigate priors and architectures, and in the long term, may demystify the dark art of architecture creation, even perhaps turning architectures into a continuous optimizable parameter of the network. Code is available at https://github.com/vsubramaniam851/untrainable-networks.
@@ 1 Introduction
When creating neural networks, we tend as a community to follow recipes that select among a few architectures known to work for particular tasks (Ren et al., 2021; Cong & Zhou, 2023; Goodfellow et al., 2014). Architecture is critical. The gains made on tasks like object recognition are attributable to imposing an inductive bias, i.e., a prior, on the design of new architectures (Goyal & Bengio, 2022; Bachmann et al., 2024). Convolutional networks unlocked many vision problems (Krizhevsky et al., 2012; He et al., 2016a) and the recent advent of the Transformer (Vaswani, 2017; Devlin, 2018; Achiam et al., 2023) did the same for language. Despite this, finding new architectures and overcoming the limitations of existing architectures remains a sort of “dark art”. While an architecture imposes some prior, we often do not fully understand what that prior is. One example of this that remains an open discussion is the role of residual connections in making very deep convolutional networks easier to train (Jastrzebski et al., 2017). If we fully understood the priors that architectures imposed, we could translate between priors and architectures freely – either by specifying the prior we want and then deriving the appropriate architecture or by dispensing with architectures and imposing the prior directly.
@@
Recent theorems (Poggio & Fraser, 2024) state that for each function which is efficiently Turing computable, there exists a deep network that can approximate it well. Furthermore, a graph representing such a function is compositionally sparse; that is, the nodes of the associated Directed Acyclic graph (DAG) represent constituent functions with a small effective dimensionality. A reasonable conjecture is that neural networks with an architecture which is similar to the DAG of the unknown target function are especially successful in learning it, as it is the case for convolutional networks for image recognition and similar tasks. Because we do not understand the relationship between the kinds of priors on the target functions that different architectures impose, even simple questions have no known answer, such as, does there exist an initialization of an FCN that makes it behave like a CNN, though the graphs of the function they represent are fundamentally different?
@@
We introduce a new tool which we term guidance to study this problem and to expand the space of viable networks. Given a target network, we guide it with a guide network. In addition to the target’s original loss, the target attempts to match the representation of its intermediate layers to those of the guide. We use a measure of representational similarity (Kornblith et al., 2019; Cristianini et al., 2001; Cortes et al., 2012), also termed a neural distance function, to compute the distance between representations of two arbitrary layers. Neural distance functions are often used in neuroscience to compare activity in networks and brains (Schrimpf et al., 2018; Conwell et al., 2021a; Subramaniam et al., 2024). In light of recent work that shows that networks of very different architectures have internal activity that is extremely similar to one another (Han et al., 2023; Conwell et al., 2021a; b), we repurpose this distance function as a means to transfer priors between networks layer by layer.
input imageconv layerbatch normreluconv layerbatch normreluconv layerbatch normrelupoolingoutput x12 Guide networkinput imagelinear layerbatch normrelulinear layerbatch normreluoutput x50 Target network
Figure 1: Guidance between two networks makes untrainable networks trainable. Given a target which cannot be trained effectively on a task, e.g., a fully connected network which immediately overfits on vision tasks, we guide it with another network. In addition to the target’s cross-entropy loss, we encourage the network to minimize the representational similarity between target and guide activations layer by layer. The guide can be untrained, i.e., randomly initialized. This procedure transfers the inductive biases from the architecture of the guide to the target. The guide is never updated. The target undergoing guidance no longer immediately overfits and can now be trained. Here we show an untrained ResNet guiding a deep fully connected network to perform object classification. The FCN alone overfits, the guided version can now be optimized. It has gone from untrainable to trainable.
@@
    We develop guidance to transfer priors between networks using representational alignment, investigate one representational alignment method, centered kernel alignment, CKA (Kornblith et al., 2019), and apply it to several classical problems.
@@
    Image classification improvements: Deep or wide fully connected networks cease overfitting when guided by a ResNet. No-skip CNNs close much of the gap with ResNets when guided by a ResNet.
@@
    Sequence modeling improvements: RNNs significantly improve their copy and paste accuracy when guided by a Transformer. Transformers increase their parity accuracy when guided by an RNN. RNNs close much of the gap to a Transformer when guided by one.
@@
    Many of these gains exist even when the guide network only transfers over its priors, not its knowledge, i.e., the guide is untrained and is never updated.
  5. 5.
    By disconnecting the guide early, we provide evidence for a better initialization regime for fully connected networks that avoids immediate overfitting.
Guidance can serve as a new tool for investigating the role of architecture and the relationship between priors and architectures. It can also help discover new initialization methods, for which no known tool exists. Guidance also has other potential applications which we leave to future work and only briefly mention in the conclusion, such as addressing some of the weaknesses of neural architecture search.
Our work has a number of limitations. We aimed for coverage of many tasks instead of maximal performance on any one task. This would have required us to carefully tune the hyperparameters involved. We preferred to show how guidance works in general rather than in cherry-picked or carefully tuned settings. To that end, we also did not optimize networks to convergence, nor did we attempt to experiment with other optimizers. Once we reproduced a well-known problematic training phenomenon, we showed that it could be overcome. We consider a network trainable enough to overcome a problem when the original problem disappears. For example, successfully training fully connected networks for object recognition was hopeless because they immediately overfit; using our guidance method they no longer do so. This does not mean that they are necessarily useful as object recognizers at present. In the case of fully connected networks, their present performance with guidance training is too low, but with additional work, we believe their performance could be substantially increased now that their train and test loss are moving in the right direction. In some cases, by applying guidance, we do see large useful improvements, such as with deep CNNs, RNNs and Transformers, although much more remains to be exploited in this domain, as well.
@@ 2 Related Work
Representational Distance: Our method builds on several metrics that measure distance between high-dimensional activations extracted from neural networks or activity in the brain (Klabunde et al., 2023). Some of these distance metrics make comparisons based on kernel matrices (Kornblith et al., 2019; Cristianini et al., 2001; Cortes et al., 2012) or relative distances (Kriegeskorte et al., 2008) between sample representations in a set. Others compute linear (Wehbe et al., 2014; Schrimpf et al., 2018) or orthogonal projections (Beauducel, 2018) from one set of representations to another. Others use canonical correlation analysis which finds linear relationships between pairs of vectors (Raghu et al., 2017; Morcos et al., 2018). These metrics are designed based on a set of desired invariant properties such as permutation invariance or invariance to linear transformations.
@@
Such approaches have been commonly applied in neuroscience for measuring representational distance of activations from networks and activity in the brain to understand which neural networks are architecturally most similar to the brain (Wehbe et al., 2014; Conwell et al., 2021a; Subramaniam et al., 2024; Goldstein et al., 2020). Under this context, Han et al. (2023) has shown the inability of current representational distance metrics to distinguish representations based on architecture. This provides a basis for our experiments to use these metrics to align representations across different neural networks. This paper gives the foundation for our intuition that networks may have similar representations that allow for transferring inductive biases from one network to another. Moreover, like Han et al. (2023), we use centered kernel alignment as our main metric of representational alignment. Recent work has explored the relationship between representational and functional distance of networks, discovering transformations between activations of networks and when these make networks functionally equivalent (Klabunde et al., 2023; Bertalan et al., 2024).
@@
Untrainable Networks: In this work, we focus on applications of FCNs and plain CNNs for image classification and RNNs for sequence modeling. Prior work has designed such approaches for these tasks, although unsuccessfully relative to the guide networks we use to improve training.
@@
FCNs have been applied for image classification, where small feed-forward networks were trained on object recognition datasets. These networks were designed with 3-5 hidden layers and less than 100 units per layer (Ma & Khorasani, 2004; Bebis & Papadourakis, 1992; Khasnobish et al., 2012; Oh & Suen, 2002). The performance of these networks was very low, where the goal of the paper was to only maximize training fit rather than generalization performance (Ma & Khorasani, 2004; Bebis & Papadourakis, 1992). Other methods were introduced to prevent overfitting in fully connected networks using topological structure (Schittenkopf et al., 1997) or early stopping (Caruana et al., 2000), but these methods had poor fits on the training set due to complex architectural design or complex hyperparameter tuning. Deep convolutional networks have been applied to image classification (Krizhevsky et al., 2012; Bengio et al., 1993) as well but face problems with vanishing gradients, preventing deep stacking of convolutional networks.
In sequence modeling, RNNs are the classical approach (Schuster & Paliwal, 1997; Pearlmutter, 1990; Connor et al., 1994; Hammer, 2000) suffering from vanishing and exploding gradients (Hochreiter, 1998). This limited early RNNs from being applicable to longer sequence lengths which was important for tasks that required memorization (Graves, 2014). Several additional tricks were developed to improve gradient flow, but ultimately new architectures like long-short term memory networks(Hochreiter, 1997) and transformers (Vaswani, 2017) led to stronger improvements. It has also been observed that transformers are untrainable on specific formal language (Bhattamishra et al., 2020) tasks that require reasoning over a full sequence instead of a sequence-to-sequence task. These tasks are successful with RNNs but not with transformers.
@@
Model Distillation: Guidance shares a resemblance with model distillation (Hinton, 2015; Gou et al., 2021; Sanh, 2019; Hsieh et al., 2023). Distillation transfers knowledge from a teacher model to a student model by introducing a new component to the loss function that enforces the student model to behave like the teacher model (Kim et al., 2021; Zhou et al., 2021). This can consist of penalizing the KL-divergence between the logit predictions of the student and teacher model. Other, more expressive variants have been proposed (Ahn et al., 2019; Chen et al., 2021b; Huang et al., 2022; Jin et al., 2023; Dai et al., 2022). In these settings, the teacher model is larger, by parameter count, and more performant. Distillation operates on model predictions rather than representations.
@@
Representation-based distillation (Tian et al., 2019; Chen et al., 2021a; Lin et al., 2020) and alignment techniques have been proposed to improve alignment between two networks. Certain works have proposed contrastive loss functions on output representations to distill teacher information into the representation space of the student. Other methods introduce correlation congruence or similarity preserving metrics for aligning two networks. Methods have been proposed that use CKA as an alignment approach between representations of two networks or with representations in the brain with notable improvement in network performance (Saha et al., 2022; Dapello et al., 2022).
@@
We distinguish guidance from distillation. Guidance can use a smaller untrained guide instead of a larger trained teacher. This is due to guidance operating over intermediate activations of the network rather than the output of the network probabilities or output features, like distillation does. Guidance also operates at many levels at the same time, aligning many layers at once. This helps address the credit assignment problem that gradient descent has when tuning weights early in a network. We also consider many more networks for guidance than is traditional for distillation including networks which have very different architectures like Transformers to RNNs. Distillation is usually carried out between two closely related architectures. We apply guidance to do the opposite.
@@ 3 Methods
Guidance introduces a term in the loss of a target network, $\mathcal{N}^{T}$, to encourage representational alignment with a guide network, $\mathcal{N}^{G}$. Only the parameters, $\theta^{T}$, of the target are updated — the guide’s parameters, $\theta^{G}$, remain fixed. Per minibatch, representational similarity is computed between the activations of $i^{G}$th layer of the guide, $\mathbf{A}^{G}_{i^{G}}(\theta^{G})$, and activations from a corresponding layer $i^{T}$ of the target, $\mathbf{A}^{T}_{i^{T}}(\theta^{T})$. We refer to the correspondence between layers of the guide $\{i^{G}\}$ and layers of the target $\{i^{T}\}$ as $I$. While this correspondence, $I$, could be complex as any two architectures can form a guide/target pair, here we choose architectures that make the correspondence obvious as is discussed later. For example, the stacked RNNs and Transformers have the same number of layers in our experiments.
The target and guide receive the same input. Per minibatch, we collect activations from intermediate layers of both networks. Layers of guide network are mapped to layers of the target network; see Figure 1. We formulate the loss in terms of minimizing the representational dissimilarity, $\bar{\mathcal{M}}$, i.e., the complement of a representational similarity metric, between guide and target activations layer by layer, summing the results. We only consider the centered kernel alignment, CKA, between the activations in this publication. Many other possibilities exist. Any representational similarity function which is differentiable could be used. Efficiency or incremental computation is much more important than it is in traditional applications since this operation happens for every minibatch.
Given $\mathcal{L}_{T}$ as the original loss of the target network, the guide network’s original loss function is irrelevant. The guide need not even have been trained on the same task or the same dataset. It need not even have been trained at all. This latter setting is what allows transferring architectural priors without transferring knowledge from the guide to the target, as there is none in a randomly initialized guide. The final loss we optimize, $\mathcal{L}$ is:
@@
This minimizes a task loss while increasing alignment between the target and guide networks given the mapping between them. The mapping may be sparse, not every layer needs to be used. This is important for guidance with transformers or stacked RNNs as will be explained later. We don’t incorporate any weight on the layer-wise similarity component of the loss. Note that the guide’s parameters, $\theta^{G}$, are constants, i.e., the guide is never updated.
3.1 Centered Kernel Alignment
To compare representations, we use a representation similarity metric, $\mathcal{M}$, which corresponds to centered kernel alignment (CKA) (Kornblith et al., 2019; Cortes et al., 2012; Cristianini et al., 2001) in our setting. We specifically consider linear CKA.
${\bm{K}}={\bm{R}}{\bm{R}}^{T},{\bm{L}}={\bm{R}}^{\prime}{\bm{R}}^{\prime T}$ (2)
$\tilde{{\bm{K}}}={\bm{H}}{\bm{K}}{\bm{H}},\tilde{{\bm{L}}}={\bm{H}}{\bm{L}}{\bm{H}}$ (3)
We compute the HSIC on the Gram matrices.
$HSIC({\bm{K}},{\bm{L}})=\text{tr}(\tilde{{\bm{K}}},\tilde{{\bm{L}}})$ (4)
Finally, we define our linear CKA metric as:
¯$\bar{\mathcal{M}}({\bm{R}},{\bm{R}}^{\prime})=1-\mathcal{M}({\bm{R}},{\bm{R}}^{\prime})$ (6)
Linear CKA ranges from 00 (identical representations) to $1$ (very different representations). Because of this, we take the complement by subtracting the linear CKA from $1$ to represent dissimilarity.
3.2 Layerwise Mapping
We design a simple method for mapping guide layers to target layers as part of providing supervision. The goal of this method is to make guide and target networks architecturally agnostic, i.e. we can supervise any target network with any guide network.
As a simple approach, we evenly spread layer computations of our guide network over our target network. For example, if we consider ResNet-18 and a 50-layer FCN, we would map every convolutional ResNet layer to every second or third linear layer of the FCN. The intuition for this approach follows from the aim of discovering the same function of the guide network using a target network. Through the design of evenly spreading layers of our ResNet-18, we are guiding the FCN to find a function similar to the guide network.
For our mapping, we consider activations from all layers with tunable weights i.e. convolutional, linear, or LSTM/RNN based layers as well as activations from layer normalizations. For multiple stacked RNNs, LSTMs, or transformers, we extract feature representations from intermediate layers in the stack as well. Using all layers is useful for guidance as it provides a strong signal to induce alignment between the target and guide networks during training. We empirically find that more layers leads to stronger results. Skipping layers based on non-linear transformations reduces memory overhead associated with storing representations per batch.

### v2

@@ Abstract
We demonstrate that architectures which traditionally are considered to be ill-suited for a task can be trained using inductive biases from another architecture. We call a network untrainable when it overfits, underfits, or converges to poor results even when tuning their hyperparameters. For example, fully connected networks overfit on object recognition while deep convolutional networks without residual connections underfit. The traditional answer is to change the architecture to impose some inductive bias, although the nature of that bias is unknown. We introduce guidance, where a guide network steers a target network using a neural distance function. The target minimizes its task loss plus a layerwise representational similarity against the frozen guide. If the guide is trained, this transfers over the architectural prior and knowledge of the guide to the target. If the guide is untrained, this transfers over only part of the architectural prior of the guide. We show that guidance prevents FCN overfitting on ImageNet, narrows the vanilla RNN–Transformer gap, boosts plain CNNs toward ResNet accuracy, and aids Transformers on RNN-favored tasks. We further identify that guidance-driven initialization alone can mitigate FCN overfitting. Our method provides a mathematical tool to investigate priors and architectures, and in the long term, could automate architecture design.
Project website at https://untrainable-networks.github.io
@@ 1 Introduction
When creating neural networks, as a community, we follow recipes that select among a few architectures that are known to work for particular tasks [63, 12, 25]. Architecture is critical, encoding essential inductive biases i.e. priors that profoundly impact their learning capabilities and performance across various tasks. Convolutional nets revolutionized vision [47, 31], and Transformers reshaped language [74, 20, 1]. Despite this, architecture design is a dark art because the precise relationship between architectures and the priors they impose is poorly understood. For example, there is discussion about exactly what the role of residual connections is [39]. This reflects a broader challenge: we rarely understand exactly what inductive biases our architectures encode. Our lack of understanding makes architecture design challenging. Given new application spaces for neural networks with rising compute costs like inference-time scaling [56], this challenge has become even more relevant.
@@
Recent theorems [59] state that for each function which is efficiently Turing computable, there exists a deep network that can approximate it well. Furthermore, a graph representing such a function is compositionally sparse, that is the nodes of the associated Directed Acyclic Graph (DAG) represent constituent functions with a small effective dimensionality. A reasonable conjecture is that neural networks with an architecture which is similar to the DAG of the unknown target function are especially successful in learning it, as it is the case for convolutional networks for image recognition and similar tasks. However, empirically testing or transferring those structural priors remains an open challenge. Because we do not understand the relationship between the kinds of priors on the target functions that different architectures impose, even simple questions have no known answer. For example, can an FCN’s initialization be tailored to mimic a CNN’s inductive bias, despite their distinct graphs?
@@
To bridge this gap, we introduce a novel empirical tool, guidance. Given a target network, we guide it with a guide network. In addition to the target’s original loss, the target attempts to match the representation of its intermediate layers to those of the guide. We use a measure of representational similarity [45, 17, 16], also termed a neural distance function, to compute the distance between representations of two arbitrary layers. Neural distance functions are often used in neuroscience to compare activity in networks and brains [68, 14, 71]. In light of recent work that shows that networks of very different architectures have internal activity that is extremely similar to one another [30, 14, 15], we repurpose this distance function as a means to transfer priors between networks layer by layer. Surprisingly, even a randomly initialized guide—incapable of solving the task—yields large performance gains, proving architectures alone encode powerful priors. This surprising finding demonstrates that neural architectures alone, independent of parameter training, impose meaningful inductive biases that are useful for downstream tasks.
Figure 1: Guidance makes untrainable networks trainable via representational similarity. Given a target which cannot be trained effectively on a task, we train this target with a layerwise representational‐alignment term against a fixed guide—trained or random—which remains unchanged during task training. This transfers only the guide’s architectural bias, turning a network that would otherwise overfit or underfit into one that learns effectively (e.g., a deep FCN guided by a random ResNet for image classification).
@@
    We develop guidance to transfer priors between networks using representational alignment and investigate one representational alignment method, centered kernel alignment, CKA [45].
@@
    We empirically differentiate between architectural and trained inductive biases, showing that architectural priors alone can significantly improve network performance. This underscores the intrinsic structural power of architectures independent of learned parameters.
@@
    We show that RNNs significantly improve their copy-and-paste accuracy when guided by a Transformer. Transformers increase their parity accuracy when guided by an RNN. RNNs close much of the gap to a Transformer on language modeling when guided by one.
@@
    We show that deep or wide fully connected networks stop overfitting when guided by a ResNet. No-skip CNNs close much of the gap with ResNets when guided by a ResNet. Fully-connected networks stop overfitting with guidance-only initialization schemes.
Our method provides a powerful empirical tool to further theoretical insights in architectural design. Guidance enables systematic investigation of the structural foundations of successful architectures and clarifies the distinction between architectural and trained inductive biases. Minimizing CKA lets you clamp any subset of a target network’s activations onto those of a frozen guide network, sweep that clamp across architectures and priors, and observe what fails or succeeds. Guidance is a knob you can turn to inject architectural priors at will, something one cannot do with cross-entropy or weight-decay alone.
Our work has a number of limitations. We aimed for coverage of many tasks instead of maximal performance on any one task. This would have required us to carefully tune the hyperparameters involved. We preferred to show how guidance works in general rather than in cherry picked and carefully tuned settings. To that end, we also did not optimize networks to convergence, nor did we attempt to experiment with other optimizers. Once we reproduced a well-known problematic training phenomenon, we showed that it could be overcome. We consider a network trainable and a problem to be overcome when the original problem disappears. For example, successfully training fully connected networks for object recognition was hopeless because they immediately overfit; using our guidance method they no longer do so. This does not mean that they are necessarily useful as object recognizers at present. In the case of fully connected networks, their present performance with guidance training is too low, but with additional work we believe their performance could be substantially increased now that their train and test loss are moving in the right direction. In some cases, by applying guidance, we do see large useful improvements, such as with RNNs and Transformers, as well as deep CNNs, although much more remains to be exploited there too.
@@ 2 Related Work
Representational Distance: Our method builds on several metrics that measure distance between high-dimensional activations extracted from neural networks or activity in the brain [44]. Some of these distance metrics make comparisons based on kernel matrices [45, 17, 16] or relative distances [46, 55] between sample representations in a set. Others compute linear [75, 68] or orthogonal projections [5] from one set of representations to another. These metrics are designed based on a set of desired invariant properties such as permutation invariance or invariance to linear transformations.
@@
Such approaches have been commonly applied in neuroscience for measuring representational distance of activations from networks and activity in the brain to understand which neural networks are architecturally most similar to the brain [75, 14, 71, 24]. Under this context, Han et al. [30] has shown the inability of current representational distance metrics – specifically the metric used here, centered kernel alignment – to distinguish representations based on architecture. This paper provides the foundation for our intuition that networks may have similar representations that allow for transferring inductive biases from one network to another.
@@
Untrainable Networks: This work examines RNNs and transformers for sequence modeling and FCNs and plain CNNs for image classification. Prior work explored similar approaches but performed poorly compared to the guide networks we leverage for improved training.
In sequence modeling, classical RNNs [69, 58, 13, 29] were constrained by vanishing and exploding gradients [35], making them unsuitable for long sequence tasks requiring memorization [27]. Gradient flow techniques were developed, but significant progress came from architectures like LSTMs [34] and transformers [74]. Transformers, however, have been found untrainable on formal language tasks requiring full-sequence reasoning, where RNNs succeed [8].
@@
For image classification, small feed-forward networks with 3-5 hidden layers and fewer than 100 units per layer were trained on object recognition datasets [53, 6, 41, 57]. These efforts prioritized training fit over generalization performance and achieved low results [53, 6]. Strategies to reduce overfitting, such as topological structure [67] or early stopping [9], were hindered by complex designs and hyperparameter tuning, leading to poor training fits. Further methods used alignment between thin deep FCNs and wide shallow FCNs to prevent overfitting, an approach similar to our paper [64]. Deep convolutional networks were also applied to image classification [47, 7] but struggled with vanishing gradients, limiting their depth.
@@
Model Distillation: Guidance shares a resemblance with model distillation [33, 26, 66, 36]. Distillation transfers knowledge from a teacher model to a student model by introducing a new component to the loss function that enforces the student model to behave like the teacher model [42, 76]. This usually consists of penalizing the KL-divergence between the logit predictions of the student and teacher model.
@@
Representation-based distillation [72, 10, 50] and alignment techniques have been proposed to improve alignment between two networks. Certain works have proposed correlation congruence or similarity preserving metrics [62] for aligning two networks, particularly as a way to do architecture search between CNNs [4]. Methods have been proposed that use CKA as an alignment approach between representations of two networks or with representations in the brain with notable improvement in network performance [65, 18].
@@
We distinguish guidance from distillation. Guidance can use a smaller untrained guide instead of a larger trained teacher. This is due to guidance operating over intermediate activations of the network instead the output of the network probabilities or output features, like distillation does. Guidance also operates at many levels at the same time, aligning many layers at once. This helps address the credit assignment problem that gradient descent has when tuning weights early in a network. We also consider many more networks for guidance than is traditional for distillation including networks which have very different architectures like Transformers to RNNs. Distillation is usually carried out between two closely related architectures. We apply guidance to do the opposite.
@@ 3 Methods
Guidance introduces a term in the loss of a target network, $\mathcal{N}^{T}$, to encourage representational alignment with a guide network, $\mathcal{N}^{G}$. We update only the target’s parameters, $\theta^{T}$, while keeping the guide’s parameters, $\theta^{G}$, frozen. On each minibatch, we compute a similarity metric $\mathcal{M}$ (e.g., CKA) between guide‐layer activations $\mathbf{A}^{G}_{i^{G}}$ and matched target activations $\mathbf{A}^{T}_{i^{T}}$. We refer to the correspondence between layers of the guide $\{i^{G}\}$ and layers of the target $\{i^{T}\}$ as $I$. While this correspondence, $I$, could be complex as any two architectures can form a guide/target pair, here we choose architectures that make the correspondence obvious as is discussed later. For example, the stacked RNNs and Transformers have the same number of layers in our experiments.
The target and guide receive the same input. Per minibatch, we collect activations from intermediate layers of both networks. Layers of guide network are mapped to layers of the target network; see fig.˜1. We formulate the loss in terms of minimizing the representational dissimilarity, $\bar{\mathcal{M}}$, i.e., the complement of a representational similarity metric, between guide and target activations layer by layer, summing the results. Here we use linear CKA, though any differentiable similarity metric could plug into Eq. (1). Efficiency or incremental computation is much more important than it is in traditional applications since this operation happens for every minibatch. We discuss details in section˜B.1.
Given $\mathcal{L}_{T}$ as the original loss of the target network, the guide network’s original loss function is irrelevant. The guide could be pretrained on another task. In fact, it need not even have been trained at all, only its architecture shapes the target. This latter setting is what allows transferring architectural priors without transferring knowledge from the guide to the target, as there is none in a randomly initialized guide. See eq.˜1 for an overall loss. We discuss details in appendix˜B.
@@
Equation˜1 minimizes a task loss while increasing alignment between the target and guide networks given the mapping between them. The mapping may be sparse; not every layer needs to be used. This is important for guidance with transformers or stacked RNNs, as will be explained later. Note that the guide’s parameters, $\theta^{G}$, are constants, i.e., the guide is never updated.
Metrics like CKA can capture and encode inductive biases in neural network computations. For instance, CKA is a measure that depends on second-order statistics, specifically pairwise sample distance matrices. Architectural choices imprint distinct features on those statistics. For example, consider local receptive fields in a convolutional layer. Units that cover neighboring pixels receive correlated input, and this is reflected in our activations. Such correlations are reflected in our distance matrices, and these can be transferred to distance matrices associated with FCN layers that lack local correlations. Similarly, weight sharing, where the same kernel is applied at every spatial location, will also be reflected in a distance matrix.
Layerwise Mapping
We design a simple method for mapping guide layers to target layers as part of providing supervision. The goal of this method is to make guide and target networks architecturally agnostic i.e. we can supervise any target network with any guide network.
As a simple approach, we evenly spread layer computations of our guide network over our target network. For example, if we consider ResNet-18 and a 50-layer FCN, we would map every convolutional ResNet layer to every second or third linear layer of the FCN. Intuitively, evenly spacing guide‐to‐target matches encourages the target to approximate the guide’s compositional functions. Through the design of evenly spreading layers of our ResNet-18, we are guiding the FCN to find a function similar to the guide network.
For our mapping, we consider activations from all tunable-weight layers (convolutional, linear, or RNN/LSTM). For multiple stacked RNNs, LSTMs, or transformers, we extract feature representations from intermediate layers in the stack as well. Using all layers is useful for guidance as it provides a strong signal to induce alignment between the target and guide networks during training. We empirically find that more layers leads to stronger results. Skipping layers based on non-linear transformations reduces memory overhead associated with storing representations per batch.